OpenStudy (anonymous):

Assume that there are 12 girls and 8 boys in the neighborhood club, and a team of 6 is to be selected. (1) How many different teams can be selected? (2) How many different teams can be selected if each team must contain exactly 4 girls and 2 boys? (3) How many different teams can be selected if each team must contain both boys and girls?

OpenStudy (anonymous):

(1)three different teams(2)three different teams(3)three different teams.

OpenStudy (dumbcow):

1) 20C6 = 38,760 2) 12C4 * 8C2 = 495*28 = 13,860 3) 20C6 - 12C6 -8C6 = 38,760 - 924 -28 = 37,808